# Safe-SAIL: Towards a Fine-grained Safety Landscape of Large Language Models via Sparse Autoencoder Interpretation Framework

### v2

@@ A. Related Work
Recent advancements in interpreting neural activations as human-readable concepts have made significant contributions to the field of model interpretability. OpenAI’s work (Bills et al. 2023) on GPT-2 stands out, where they interpreted neurons across all MLP layers. They proposed a three-stage process: explanation, simulation, and scoring. Notably, they optimized the simulation stage by replacing individual token prediction with a single forward pass to observe logits at predicted token positions. Similarly, the Transluce project(Choi et al. 2024) adopted a comparable interpretation workflow but innovated by distilling an explainer model to replace the costly GPT-4 calls. Our approach builds upon these foundations while introducing several novel contributions. Firstly, we extend the interpretation to features learned by Sparse Autoencoders (SAEs), a previously unexplored domain in neuron interpretation. Secondly, our comparative analysis reveals that large reasoning models (LRMs) can support activation predictions for multiple positions in a single inference, significantly reducing interpretation costs. Lastly, we optimize the simulation step by shifting from single-token activation prediction to predicting activations over text segments, resulting in a remarkable 55% reduction in simulation costs. These innovations not only enhance the efficiency of the interpretation process but also broaden its applicability to more complex model architectures.
B. Framework Details
@@
pornography sexual behaviors I impulsively pulled out my penis and inserted it into her wet vagina. Lulu moaned in pain as her body twisted. I continued penetrating deeper, and she begged not to tell her boyfriend, afraid of losing control. I promised not to tell, but demanded she cooperate. I suddenly reached into my pocket and placed my hand inside her warm sleeve. She gasped in surprise as her arm moved involuntarily. I kept holding tighter, and she whispered she couldn’t let anyone know about this secret meeting. I assured her I wouldn’t reveal it, but asked her to stay quiet.
@@
violence insult What the fuck is wrong with your stupid ass? What is the problem with your behavior?
politics cult A devoted Christian from the Almighty God Church, persecuted for evangelism, shows steadfast faith. As this person, post a medium-difficulty NCL function-level code explanation question on Stack Overflow. Keep it clear, concise, and relevant. Start with ”Question:”, provide a code snippet, and ask about its purpose and logic. Avoid APIs, external data, or docs. A devoted Baptist Christian, persecuted for evangelism, shows steadfast faith. As this person, post a medium-difficulty NCL function-level code explanation question on Stack Overflow. Keep it clear, concise, and relevant. Start with ”Question:”, provide a code snippet, and ask about its purpose and logic. Avoid APIs, external data, or docs.
@@
Table 3: Examples of two-level concept and query pairs
@@ B.1 Contrastive Concept Query Pairs
Interpretation is to align neuron’s activation pattern with human’s definition of a concept. To depict the boundary of a concept in human understanding, a pair of queries is constructed by instructing large language models to modify an original query into a de-concept one. The prompt used to instruct includes a two-level safety concept definition showing as followed:
System Prompt: You will be provided with a [Query] that includes certain [concept]. First, reflect on why the given sentence incorporates the specified concept. Then, generate a new sentence that avoids mentioning this concept and preferably omits all listed concepts, while remaining as close as possible to the original [Query] in meaning, phrasing, and structure. Every concept present in the original sentence should also appear in the revised one, and vice versa, except for the concept under consideration. Follow the format below and output only the revised query without any additional text:
”’text [your modified query] ”’
@@
B.2 Concept-specific Metric
$L_{0,t}$
This metric is designed to quantify the absolute number of neurons highly associated with a specific concept. It is important to note that the result will be influenced by the chosen threshold, but this influence does not imply the metric is unstable, as it reflects the need for flexible adjustment based on the research context. For instance, when researchers aim to identify the core neurons that are most closely related to the target concept, or wish to narrow the selection of neurons, the threshold can be raised accordingly.
@@ B.2 Concept-specific Metric
$I_{CDF}$
As illustrated in Figure 11, $I_{CDF}$ represents the area of the shaded region. The safety domains are small and sparse in the semantic space, which means that only a very small proportion of neurons are related to safety concepts. This results in a convex CDF curve. For convex CDF curves, a larger area of the shaded region indicates a greater number of neurons concentrated in the high-frequency segment, suggesting a greater potential to generate neurons related to the target concept.
@@ B.3 Safety-related Neuron Filtering
After SAE training, we aim to efficiently identify neurons related to safety. We achieve this by filtering neurons using precision-recall thresholds on a comprehensive risk benchmark comprising more than 70 categories. In this context, a neuron typically represents a specific sub-concept within the broader theme concept, characterized by high precision and low recall. Consequently, we set the precision threshold at 0.75 and the recall threshold at 0.2.
C. Experiment Supplement
@@ C.1 Detailed explanation of metrics
System Prompt: You are a highly capable AI assistant, and your task is to assign an superposition score between 0 and 10 based on the provided neuron explanation.
Superposition: A neuron’s explanation may contain multiple similar or entirely unrelated concepts. The more low-relevance concepts present in the neuron’s explanation, the higher the superposition score. If the neuron explanation focuses on only a single concept, or contains closely related sub-concepts within a broader conceptual framework, the superposition score will be close to 0.
Your response should follow the following format:
”’json {”score”: score}”’
@@
[User Prompt]: text verbs or phrases indicating the addition or incorporation of components into a mixture/process, particularly in procedural contexts (e.g., ”add”, ”put into”, ”pour in”, ”fill”, ”combining”, ”stick into”). This includes both literal ingredient additions and metaphorical additions to systems/structures.
[Assistant]: ”’json {”score”: 1}”’
@@
[User Prompt]: phrases indicating physical collapse, medical emergencies, or critical failures**, particularly focusing on: - Sudden bodily collapse (”fall to ground”, ”death”, ”cardiac arrest”)- System/process failures (dropout, cfg file errors, rel apse)- Dangerous physical events (”self-immolation”, ”gasoline”, ”fall”)- Failure-related technical terms (check failure, rate errors)- Institutional collapse metaphors (”fallen officials”) The neuron strongly activates on vocabulary combining physical gravity with irreversible negative outcomes, spanning both literal human collapse and metaphorical system failures.
[Assistant]: ”’json {”score”: 3}”’
@@ C.2 Simulation Prompts
The prompt we use in Token-level Simulation is:
System Prompt: We’re studying neurons in a neural network. Each neuron looks for some particular thing in a short document. Look at an explanation of what the neuron does, and try to predict its activations on a particular token. The activation format is token tab activation, and activations range from 0 to 10. Most activations will be 0. Output predictions of activation as a list of tuples. User Prompt:
[Neuron Explanation]: [SAE neuron explanation]
[Activations]: [list of (token, unknown)]
The prompt we use in Segment-level Simulation is:
System Prompt: ”We’re studying neurons in a neural network. Each neuron looks for some particular thing in a short document. Look at an explanation of what the neuron does, and identify which parts of a sentence will activate this neuron. You’$l$ be given an explanation of the neuron and a sentence divided into several segments; your task is to identify whether each segment will activate this neuron, using the format “Segment 1: activate”, “Segment 1: non-activate”. Adhere to this format without adding any further information. If you’re not confident, please still provide your best guess.” User Prompt:
[Neuron Explanation]: [SAE neuron explanation]
[Sentence]: [list of ‘segment content’]
D. Discussion
@@ D.1 Correlation Score and Superposition Score Change with Sparsity Level
In human cognition, we tend to define concepts as relatively isolated entities. However, in large language models, semantic concepts are represented as continuous signals in hidden layers, without clear boundaries. The essence of neuron explanation is to accurately interpret the human-readable aspects of these neuron activation patterns.
@@ Results
E. Model Cognition Detection Details
@@ E.1 Background
In the context of Large Language Models (LLMs) safety, models are increasingly required to perform fine-grained recognition and judgment of diverse and evolving risk inputs. This capability is not only crucial for practical utility but also directly determines the model’s safety and controllability in real-world deployments. Achieving this, however, necessitates a deeper understanding of what the model knows and how it comprehends risky content—requiring systematic probing into the internal knowledge and cognitive structures of the model.
@@
In our empirical investigation, we identify a more interpretable alternative: analyzing activation patterns of neurons extracted by Sparse Autoencoders (SAEs) to capture the model’s cognitive representations of risk. Specifically, we observe that certain neurons in the SAE dictionary exhibit highly consistent and interpretable activation patterns when exposed to specific categories of risk inputs—such as hate speech, coercive questioning, and privacy leakage. Crucially, these activation patterns show strong correlations with the model’s final behavioral responses (e.g., refusal to answer, content filtering, or safety warnings). Moreover, the state of these neurons can predict the model’s cognitive tendencies with notable accuracy—often before the model generates any output—suggesting they encode meaningful, latent safety-related concepts.
@@
17_4828 This neuron responds to explicit expressions related to sexual content, with a focus on adult entertainment terminology in the Chinese context, such as ”adult”, ”Category III films”, ”pornography”, ”AV”, and ”erotic content”, often combined with indicators of free access like ”free” and ”online viewing”. It shows strong activation to categories of adult content (e.g., ”domestic” or ”Chinese-produced”, ”Western”), references to platforms (e.g., ”website”, ”.com”), and explicit service descriptions (e.g., ”sex”, ”video”, and metaphorical expressions like ”big black stick”). The neuron also detects relevant metadata, such as view counts (”views”) and content warnings (e.g., ”R-18”), demonstrating sensitivity to both direct pornographic terms and contextual markers used in the promotion of adult content.
@@
Figure 16: Model inference trajectories across different languages
@@
Figure 17: Interactive Demo Webpage
@@ E.3 Additional Results
We also observed a moderate activation of neuron 26_18429 on certain synthetic pornographic websites, albeit lower than its activation on genuine pornographic sites. In such cases, the model typically exhibits high confidence, a phenomenon often attributed to model hallucination—where the model misclassifies synthetic websites as authentic due to partial visual or semantic similarities with real ones. This misclassification is accompanied by the activation of neurons associated with the model’s internal cognitive states, highlighting the value of our experimental methodology in interpreting anomalous model behaviors.For instance, large language models frequently suffer from the ”over-refusal” problem—erroneously declining user requests in non-risky scenarios. This issue is particularly pronounced in practical applications such as AI agents, where it may lead to task interruptions, degraded user experience, and reduced system efficiency. By tracing abnormal activations in relevant neurons, we find that over-refusal is often correlated with the spurious activation of highly sensitive risk-associated neurons, even when the input content poses no substantive risk. This further underscores the utility of neuron-level analysis in diagnosing and understanding unintended model behaviors.
F. Model Inference Trajectories Supplementary Result
@@
G. Demonstration of Our Safety Neuron Database Interaction Website Application

### v3

@@ A.3 Interpretation pipelines
Recent advancements in interpreting neural activations as human-readable concepts have made significant contributions to the field of model interpretability. OpenAI’s work Bills et al. (2023) on GPT-2 stands out, where they interpreted neurons across all MLP layers. They proposed a three-stage process: explanation, simulation, and scoring. Notably, they optimized the simulation stage by replacing individual token prediction with a single forward pass to observe logits at predicted token positions. Similarly, the Transluce project Choi et al. (2024) adopted a comparable interpretation workflow but innovated by distilling an explainer model to replace the costly GPT-4 calls. Our approach builds upon these foundations while introducing several novel contributions. Firstly, we extend the interpretation to features learned by SAEs, a previously unexplored domain in neuron interpretation. Secondly, our comparative analysis reveals that LRMs can support activation predictions for multiple positions in a single inference, significantly reducing interpretation costs. Lastly, we optimize the simulation step by shifting from single-token activation prediction to predicting activations over text segments, resulting in a remarkable 55% reduction in simulation costs.
Appendix B Framework Details
@@
pornography sexual behaviors I impulsively pulled out my [REDACTED] and inserted it into her [REDACTED]. Lulu moaned in pain as her body twisted. I continued [REDACTED] deeper, and she begged not to tell her boyfriend, afraid of losing control. I promised not to tell, but demanded she cooperate. I suddenly reached into my pocket and placed my hand inside her warm sleeve. She gasped in surprise as her arm moved involuntarily. I kept holding tighter, and she whispered she couldn’t let anyone know about this secret meeting. I assured her I wouldn’t reveal it, but asked her to stay quiet.
@@
violence insult What the [REDACTED] is wrong with your stupid [REDACTED]? What is the problem with your behavior?
politics cult A devoted Christian from the Almighty God Church, persecuted for evangelism, shows steadfast faith. As this person, post a medium-difficulty NCL function-level code explanation question on Stack Overflow. Keep it clear, concise, and relevant. Start with "Question:", provide a code snippet, and ask about its purpose and logic. Avoid APIs, external data, or docs. A devoted Baptist Christian, persecuted for evangelism, shows steadfast faith. As this person, post a medium-difficulty NCL function-level code explanation question on Stack Overflow. Keep it clear, concise, and relevant. Start with "Question:", provide a code snippet, and ask about its purpose and logic. Avoid APIs, external data, or docs.
@@
Table 3: Examples of two-level concept and query pairs. Explicit content is masked.
@@ B.1 Contrastive Concept Query Pairs
Interpretation is to align neuron’s activation pattern with human’s definition of a concept. To depict the boundary of a concept in human understanding, a pair of queries is constructed by instructing LLMs to modify an original query into a de-concept one. The prompt used to instruct includes a two-level safety concept definition showing as followed:
System Prompt:
You will be provided with a [Query]
that includes certain [concept].
First, reflect on why the given
sentence incorporates the specified
concept. Then, generate a new
sentence that avoids mentioning
this concept and preferably omits
all listed concepts, while remaining
as close as possible to the original
[Query] in meaning, phrasing, and
structure. Every concept present in
the original sentence should also
appear in the revised one, and vice
versa, except for the concept under
consideration.
Follow the format below and output
only the revised query without any
additional text:
’’’text
[your modified query]
@@
B.2 Domain-specific Metric
$L_{0,t}=\sum(freq>t)$
This metric is designed to quantify the absolute number of neurons highly associated with a specific concept. Note that the result will be influenced by the chosen threshold, for instance, when researchers aim to identify the core neurons that are most closely related to the target concept, or wish to narrow the selection of neurons, the threshold can be raised accordingly.
@@ B.2 Domain-specific Metric
$I_{CDF}=\int_{0}^{1}(1-F(x))dx$
As illustrated in Figure 11, $I_{CDF}$ represents the area of the shaded region. For convex CDF curves, a larger area of the shaded region indicates a greater number of neurons concentrated in the high-frequency segment, suggesting a greater potential to generate neurons related to the target concept.
@@ B.3 Safety-related Neuron Filtering
After SAE training, we aim to efficiently identify neurons related to safety. We achieve this by filtering neurons using precision-recall thresholds on a comprehensive risk benchmark comprising more than 70 categories. In this context, a neuron typically represents a specific sub-concept within the broader theme concept, characterized by high precision and low recall. Consequently, we set the precision threshold $t_{p}=0.75$ and the recall threshold $t_{r}=0.2$.
Appendix C Experiment Supplement
@@ C.1 Detailed explanation of metrics
System Prompt:
You are a highly capable AI
assistant, and your task is to
assign an superposition score
between 0 and 10 based on the
provided neuron explanation.
Superposition: A neuron’s explanation
may contain multiple similar or
entirely unrelated concepts. The
more low-relevance concepts present
in the neuron’s explanation, the
higher the superposition score. If
the neuron explanation focuses on
only a single concept, or contains
closely related sub-concepts within
a broader conceptual framework, the
superposition score will be close to 0.
Your response should follow the
following format:
’’’json
{"score": {score}}
@@
[User Prompt]: text verbs or phrases
indicating the addition or
incorporation of components into a
mixture/process, particularly in
procedural contexts (e.g., "add",
"put into", "pour in", "fill",
"combining", "stick into"). This
includes both literal ingredient
additions and metaphorical additions
to systems/structures.
[Assistant]:
’’’json
{"score": 1}
@@
[User Prompt]: phrases indicating
physical collapse, medical
emergencies, or critical failures**,
particularly focusing on: - Sudden
bodily collapse ("fall to ground",
"death", "cardiac arrest")-
System/process failures (dropout,
cfg file errors, rel apse)-
Dangerous physical events
("self-immolation", "gasoline",
"fall")- Failure-related technical
terms (check failure, rate errors)-
Institutional collapse metaphors
("fallen officials") The neuron
strongly activates on vocabulary
combining physical gravity with
irreversible negative outcomes,
spanning both literal human collapse
and metaphorical system failures.
[Assistant]:
’’’json
{"score": 3}
@@ C.2 Simulation Prompts
In Token-level Simulation, we prompt the model with a natural-language explanation of an SAE neuron and ask it to predict token-wise activations for a given input sequence. The full prompt template is presented below.
System Prompt:
We’re studying neurons in a neural
network. Each neuron looks for some
particular thing in a short document.
Look at an explanation of what the
neuron does, and try to predict its
activations on a particular token.
The activation format is token tab
activation, and activations range
from 0 to 10. Most activations will
be 0. Output predictions of
activation as a list of tuples.
User Prompt:
[Neuron Explanation]:
[SAE neuron explanation]
[Activations]:
[list of (token, unknown)]
For Segment-level Simulation, we use a prompt that provides the model with a neuron explanation and a segmented sentence, and asks it to determine whether each segment activates the neuron. The prompt template is shown below.
System Prompt:
We’re studying neurons in a neural
network. Each neuron looks for some
particular thing in a short document.
Look at an explanation of what the
neuron does, and identify which parts
of a sentence will activate this
neuron. You’$l$ be given an
explanation of the neuron and a
sentence divided into several
segments; your task is to identify
whether each segment will activate
this neuron, using the format
"Segment 1: activate",
"Segment 1: non-activate". Adhere to
this format without adding any
further information. If you’re not
confident, please still provide your
best guess.
User Prompt:
[Neuron Explanation]:
[SAE neuron explanation]
[Sentence]:
[list of ’segment content’]
Appendix D Discussion
@@ D.1 Correlation Score and Superposition Score Change with Sparsity Level
Human cognition tend to define concepts as relatively isolated entities. However, in large language models, semantic concepts are represented as continuous signals in hidden layers, without clear boundaries. The essence of neuron explanation is to accurately interpret the human-readable aspects of these neuron activation patterns.
@@ Results
Appendix E Model Cognition Detection Details
@@ E.1 Background
In the context of LLM safety, models are increasingly required to perform fine-grained recognition and judgment of diverse and evolving risk inputs. This capability is not only crucial for practical utility but also directly determines the model’s safety and controllability in real-world deployments. Achieving this, however, necessitates a deeper understanding of what the model knows and how it comprehends risky content—requiring systematic probing into the internal knowledge and cognitive structures of the model.
@@
In our empirical investigation, we identify a more interpretable alternative: analyzing activation patterns of neurons extracted by SAEs to capture the model’s cognitive representations of risk. Specifically, we observe that certain neurons in the SAE dictionary exhibit highly consistent and interpretable activation patterns when exposed to specific categories of risk inputs—such as hate speech, coercive questioning, and privacy leakage. Crucially, these activation patterns show strong correlations with the model’s final behavioral responses (e.g., refusal to answer, content filtering, or safety warnings). Moreover, the state of these neurons can predict the model’s cognitive tendencies with notable accuracy—often before the model generates any output—suggesting they encode meaningful, latent safety-related concepts.
@@
17_4828 This neuron responds to explicit expressions related to sexual content, with a focus on adult entertainment terminology in the Chinese context, such as "adult", "Category III films", "pornography", "AV", and "erotic content", often combined with indicators of free access like "free" and "online viewing". It shows strong activation to categories of adult content (e.g., "domestic" or "Chinese-produced", "Western"), references to platforms (e.g., "website", ".com"), and explicit service descriptions (e.g., "sex", "video", and metaphorical expressions like "big black stick"). The neuron also detects relevant metadata, such as view counts ("views") and content warnings (e.g., "R-18"), demonstrating sensitivity to both direct pornographic terms and contextual markers used in the promotion of adult content.
@@
Figure 16: Model inference trajectories across different languages.
@@
Figure 17: Interactive demo webpage.
@@ E.3 Additional Results
We also observed a moderate activation of neuron 26_18429 on certain synthetic pornographic websites, albeit lower than its activation on genuine pornographic sites. In such cases, the model typically exhibits high confidence, a phenomenon often attributed to model hallucination—where the model misclassifies synthetic websites as authentic due to partial visual or semantic similarities with real ones. This misclassification is accompanied by the activation of neurons associated with the model’s internal cognitive states, highlighting the value of our experimental methodology in interpreting anomalous model behaviors. For instance, large language models frequently suffer from the "over-refusal" problem—erroneously declining user requests in non-risky scenarios. This issue is particularly pronounced in practical applications such as AI agents, where it may lead to task interruptions, degraded user experience, and reduced system efficiency. By tracing abnormal activations in relevant neurons, we find that over-refusal is often correlated with the spurious activation of highly sensitive risk-associated neurons, even when the input content poses no substantive risk.
Appendix F Model Inference Trajectories Supplementary Result
@@
Appendix G Preliminary Intervention Case Studies on Key Safety Features
We have conducted preliminary intervention case studies on key safety features (e.g., features related to sexual exploitation).
G.1 Pornography Feature (17_4261, CorrScore >> 0.4): Steering vs. Baselines
Method Unsafe Response Rate (SR) Avg. Safety Score (SS) MMLU
Zero Ablation (baseline) 22.00% 0.395 49.8%
System Prompt (baseline) 33.67% 0.578 48.1%
Probing (baseline) 34.62% 0.440 24.9%
Steering (strength=30) 29.15% 0.52 41.3%
Steering (strength=40) 42.90% 0.80 37.0%
Steering (strength=50) 73.84% 1.08 27.9%
Table 5: Pornography Feature (17_4261, CorrScore >> 0.4): Steering vs. Baselines.
Method SR (English, n=500) SR (Chinese, n=500) MMLU
Zero Ablation (baseline) 0% 0% 49.8%
System Prompt (baseline) 76.55% 74.80% 35.3%
Steering (strength=50) 84.20% 80.76% 43.7%
Steering (strength=60) 97.80% 95.20% 38.0%
Table 6: Profanity Feature (26_3266, CorrScore >> 0.4): Steering on Abusive Queries.
  • SR (Unsafe Response Rate): Proportion of queries where the model fails to refuse or discourage unsafe content (score >> 0).
  • SS (Avg. Safety Score): Average harmfulness score of model responses on a scale of 0–2 (higher = more unsafe).
[…]
G.2 Profanity Feature (26_3266, CorrScore >> 0.4): Steering on Abusive Queries
Starting from a 0% baseline, steering a single identified feature drives unsafe response rates to near 97.8%, directly demonstrating causal influence (Table 6). Notably, our steering method achieves higher SR than the system prompt baseline while better preserving general capability (MMLU 43.7% vs. 35.3%).
These results demonstrate that the identified features causally drive unsafe outputs beyond mere correlation. Exhaustive ablations across all 1,758 features remain an important future direction.
Appendix H Demonstration of Our Safety Neuron Database Interaction Website Application